OpenStudy (anonymous):

Find the quotient of the quantity negative 40 times x to the 3rd power minus 12 times x to the 2nd power plus 16 times x all over negative 4 times x. 10x2 + 3x − 4 −40x3 − 12x2 − 4 −10x2 + 3x − 4 10x2 − 3x + 4

OpenStudy (anonymous):

PLEASE HELP ME

OpenStudy (kainui):

Alright so let's take this one step at a time. I'm here to help if you mess up, so don't worry about making mistakes. Just write the symbol for each word they say. So they start out with: "negative 40 times x to the 3rd power" Write it out one step at a time, and I'll check your answer. Try to go further than that if you think you can, just don't let it overwhelm you.

OpenStudy (anonymous):

−40x3 − 12x2 − 4

OpenStudy (kainui):

Where did you get -4 from?

OpenStudy (anonymous):

Because its says negative

OpenStudy (kainui):

"40 times x to the 3rd power minus 12 times x to the 2nd power plus 16 times x" doesn't have a 4 in it.

OpenStudy (anonymous):

i cant get it

OpenStudy (kainui):

Just turn this sentence into all symbols. When you see it say "plus" you write +.. When you see "times" just write *. I know you can do it, just gotta copy each word one at a time into the symbol you already know it is. =D "40 times x to the 3rd power minus 12 times x to the 2nd power plus 16 times x"

OpenStudy (anonymous):

40*x^3-12*x^2+16*x

OpenStudy (kainui):

Perfect! Now you just need to finish it, they said it was a quotient (fancy way of calling it a fraction) "all over negative 4 times x" After that I can help you simplify it if you need it. =)

OpenStudy (anonymous):

-4*x

OpenStudy (kainui):

\[\frac{ -40x^3-12x^2+16x }{ -4x }\] Now, can you divide -4x out of the top of the fraction?

OpenStudy (anonymous):

well i cant divide if i cant find my calculater

OpenStudy (anonymous):

can u please just divide it for me please

OpenStudy (anonymous):

its b right

OpenStudy (kainui):

You need to factor out (-4x) from the top of the fraction.

OpenStudy (anonymous):

i know but i dont have a calculater

OpenStudy (debbieg):

NO, you don't need a calculator! This is algebra and simple factoring. Don't be calculator-dependent! It won't help you get any better understanding of this basic math if we just tell you the answer. You have to FACTOR the -4x out of each term in the numerator, then you can cancel it with the -4x in the denominator. So you are looking for the expression (which will be a polynomial) that goes in the "?" so that:\[-4x(?)=−40x^3−12x^2+16x\] Break it up and do it one term at a time:\[-4x(?)=−40x^3\]\[-4x(?)=−12x^2\]\[-4x(?)=16x\]